# Long-Context Linear System Identification

### v1

@@ A.3 Supremum of the noise
Using Lemma B.12, we have ${{\left\|M_{{\boldsymbol{A}}^{\star}}\right\|}}_{\rm op}\leqslant\sqrt{p}{{\left\|{\boldsymbol{A}}^{\star}\right\|}}_{\rm op}$, directly leading to (ii).
@@
Appendix B Proof of Theorems 4.1, 4.2 and 4.3
@@
Proposition B.2.
With the definitions of Definition B.1, we have the following properties:
@@
$\begin{split}\mathcal{A}_{r,p^{\prime}}(D)&={{\left\{{\boldsymbol{A}}\in\mathbb{R}^{d\times pd}\ \Big{|}\ \|\Delta_{{\boldsymbol{A}}}\|_{\rm op}\leqslant D,\ {\rm rank}(A_{i})\leqslant r,\ A_{p^{\prime}+1}=\cdots=A_{p}=0\right\}}}\,,\\ \mathcal{S}_{r,p^{\prime}}(C,D)&={{\left\{{\boldsymbol{A}}\in\mathcal{A}(D)\ \big{|}\ \|{\boldsymbol{A}}-{\boldsymbol{A}}^{\star}_{p^{\prime}}\|_{F}^{2}\leqslant CD^{2}\eta^{2}\dfrac{p^{\prime}dr\left(\ln\tau\right)^{2}}{N(T-p^{\prime})}\right\}}}\,,\end{split}$
@@
and $\tau$ is the following term:
$\tau=e\sigma_{{\mathrm{cond}}}(L_{\star})\sqrt{p^{\prime}dNT}\sqrt{\dfrac{T}{T-p^{\prime}}}\,.$
@@
$\mathcal{G}_{r,p^{\prime}}(C,D)={{\left\{{\boldsymbol{A}}\in\mathcal{A}_{r,p}(D)\ \Big{|}\ \dfrac{\|\Delta_{{\boldsymbol{A}}}\|_{F}^{2}}{\|\Delta_{{\boldsymbol{A}}}\|_{\rm op}^{2}}\leqslant C\eta^{2}\frac{p^{\prime}dr\left(\ln\tau\right)^{2}}{N}\right\}}}\,.$
We set $\mathcal{A}_{r,p^{\prime}}=\mathcal{A}_{r,p^{\prime}}(\infty)$ and $\mathcal{G}(C)_{r,p^{\prime}}=\mathcal{G}(C,\infty)$. Lastly, we drop the subscript $r,p^{\prime}$ when the statement is valid for all $r,p^{\prime}$.
Remark B.4.
For stable systems with $\|M_{{\boldsymbol{A}}^{\star}}\|_{\rm op}<1$ and $\|M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\|_{\rm op}<1$, the factor $\eta$ is controlled by Corollary B.10. However, for marginally stable systems or explosive systems, there is no a prior good upper bound on $\eta$, implying that the misspecification results only applies to stable systems.
B.1 Theorem statement
In this subsection, we state Theorem B.5 that generalizes the statements in Theorems 4.1, 4.2 and 4.3. We give a proof that reduces Theorem B.5 to a uniform concentration result in Theorem B.6. The proof of Theorem B.6 is deferred to Section B.5. Lastly, Corollary B.7 gives a corollary that removes the constraints on the diameter of the search set.
Theorem B.5.
Let Assumptions 3.1 and 3.4 hold. Furthermore, let Assumption 3.7 for $r<d$ and Assumption 3.9 for $p^{\prime}<p$ hold. Let ${\boldsymbol{\hat{A}}}$ be the following estimator:
@@ B.1 Theorem statement
Then, for any small $\delta>0$, there exist $C(\delta)=\mathcal{O}(\ln(1/\delta))$ such that
@@
By Corollary B.13, $\mathcal{G}(C,D)\subset\mathcal{S}(C,D)$ and thus, for any random choice of ${\boldsymbol{A}}$,
($\begin{split}\mathbb{P}\left({{\left\{{\boldsymbol{A}}\in\mathcal{S}(C(\delta),D)\right\}}}\right)&\geqslant\mathbb{P}\left({{\left\{{\boldsymbol{A}}\in\mathcal{G}(C,D)\right\}}}\right)=1-\mathbb{P}\left({{\left\{{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)\right\}}}\right)\,.\\ \end{split}$
For the choice of ${\boldsymbol{\hat{A}}}$, $\mathbb{P}\left(\mathcal{E}_{{\boldsymbol{\hat{A}}}}\right)=1$ by Corollary B.15 and
($\mathbb{P}\left({{\left\{{\boldsymbol{\hat{A}}}\in\mathcal{S}(C(\delta),D)\right\}}}\right)\geqslant 1-\mathbb{P}\left({{\left\{{\boldsymbol{\hat{A}}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)\right\}}}\mid\mathcal{E}_{\boldsymbol{\hat{A}}}\right)\,.$
@@
($\begin{split}\mathbb{P}\left({{\left\{{\boldsymbol{\hat{A}}}\in\mathcal{S}(C(\delta),D)\right\}}}\right)&\geqslant 1-\mathbb{P}\left(\mathcal{E}_{\boldsymbol{\hat{A}}}\mid{{\left\{{\boldsymbol{\hat{A}}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)\right\}}}\right)\mathbb{P}\left({{\left\{{\boldsymbol{\hat{A}}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)\right\}}}\right)\\ &\geqslant 1-\mathbb{P}\left(\mathcal{E}_{\boldsymbol{\hat{A}}}\mid{{\left\{{\boldsymbol{\hat{A}}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)\right\}}}\right)\,.\end{split}$
Then, the proof is complete by applying Theorem B.6 to the right-hand side. ∎
Theorem B.6.
Let all the assumptions of Theorem B.5 hold. Then, for any small $\delta>0$, there exist a constant $C(\delta)=\mathcal{O}(\ln(1/\delta))$ such that
@@
Corollary B.7.
Let all the assumptions of Theorem B.5 hold and consider the following estimator:
${\boldsymbol{\hat{A}}}_{\rm{OLS}}=\operatorname{argmin}_{{\boldsymbol{A}}\in\mathcal{A}(\infty)}\mathcal{L}({\boldsymbol{A}})\,.$
Then, for any small $\delta>0$, there exist $C(\delta)=\mathcal{O}(\ln(1/\delta))$ such that
$\mathbb{P}\left({\boldsymbol{\hat{A}}}_{\rm{OLS}}\in\mathcal{S}(C(\delta),D)\right)\geqslant 1-\delta\,,$
given that $NT$ satisfies the following condition:
$N(T-p^{\prime})\geqslant C\eta^{2}p^{\prime 2}dr(\ln\tau)^{2}\ln\ln D\,.$ (22)
@@
$\mathcal{A}(D)^{\mathrm{o}}\subset\mathcal{G}(C,D)\,,$
i.e., the interior of $\mathcal{A}(D)$ contains $\mathcal{G}(C,D)$. We have the following relation:
$\mathcal{A}(\infty)={{\left\{{\boldsymbol{A}}^{\prime}=\alpha{\boldsymbol{A}}\mid{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D),\alpha\geqslant 1\in\mathbb{R}\right\}}}\,.$
Then, by Theorem B.6, we have
$\begin{split}&\mathbb{P}\left(\exists{\boldsymbol{A}}\in\mathcal{A}(\infty)\setminus\mathcal{G}(C(\delta),D):\|\Delta_{\boldsymbol{A}}L_{\star}E\|_{F}^{2}\leqslant 2\eta\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)\right)\\ &\quad\quad\quad=\mathbb{P}\left(\exists\alpha\geqslant 1\in\mathbb{R},{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C(\delta),D):\|\Delta_{\alpha{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}\leqslant 2\eta\operatorname{Tr}(E^{\top}\Delta_{\alpha{\boldsymbol{A}}}L_{\star}E)\right)\\ &\quad\quad\quad=\mathbb{P}\left(\exists{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C(\delta),D):\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}\leqslant 2\eta\operatorname{Tr}(E^{\top}\Delta_{{\boldsymbol{A}}}L_{\star}E)\right)\leqslant\delta\,,\end{split}$
@@
Thus, the result is complete by applying the same argument as in Theorem B.5 where $\mathcal{A}(D)$ is replaced by $\mathcal{A}(\infty)$.
We only need to provide a $D$ such that $\mathcal{A}(D)^{\mathrm{o}}\subset\mathcal{G}(C,D)$. It is easier to ensure the inclusion $\mathcal{A}(D)^{\mathrm{o}}\subset\mathcal{S}(C,D)$. For any ${\boldsymbol{A}}\in\mathcal{S}(C,D)$, we have
$\|\Delta_{{\boldsymbol{A}}}\|_{\rm op}^{2}\leqslant p^{\prime}\|{\boldsymbol{A}}\|_{\rm op}^{2}\leqslant p^{\prime}\|{\boldsymbol{A}}\|_{F}^{2}\leqslant CD^{2}\eta^{2}\dfrac{p^{\prime 2}dr\left(\ln\tau\right)^{2}}{N(T-p^{\prime})}\,,$
from Lemma B.12. Therefore, we need to find a $D$ such that
$D^{2}\geqslant CD^{2}\eta^{2}\dfrac{T}{T-p}\dfrac{p^{\prime}dr\left(\ln\tau\right)^{2}}{N}\ln\ln D\,,$ (23)
where we make the $\ln\ln D$ factor in $C$ explicit. See Section B.5 for the details why this constant is needed. Lastly, Equation 23 is satisfied for large enough $N$ and $T$ that verifies the condition in Equation 22. ∎
B.2 Technical lemmas
In this subsection, we present simple technical results on $L_{\star},M_{{\boldsymbol{A}}}$ and $\Delta_{{\boldsymbol{A}}}$ that are used in the proof of Theorem B.5.
Lemma B.8.
@@ B.2 Technical lemmas
By Weyl’s inequality for singular values on the identity $L_{\star}=M_{{\boldsymbol{A}}^{\star}}L_{\star}+I$ from Lemma B.8,
@@
This implies the desired inequalities for $\|L_{\star}\|_{\rm op}$. For the lower bound on minimal singular value, use Lemma B.8,
@@
Applying Lemma B.9,
$\eta\leqslant\left\|M_{{\boldsymbol{A}}^{\star}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right\|_{\rm op}\|L_{\star}\|_{\rm op}\leqslant\frac{1}{1-\|M_{{\boldsymbol{A}}^{\star}}\|_{\rm op}}\left(\left\|M_{{\boldsymbol{A}}^{\star}}\right\|_{\rm op}+\left\|M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right\|_{\rm op}\right)\leqslant\dfrac{2}{1-\|M_{{\boldsymbol{A}}^{\star}}\|_{\rm op}}\,.$
Lemma B.11.
@@
By Weyl’s inequality for singular values on the identity $L_{\star}=I+M_{{\boldsymbol{A}}^{\star}}+\cdot+M_{{\boldsymbol{A}}^{\star}}^{T-1}$ from Lemma B.8,
@@
For the lower bound on minimal singular value, use Lemma B.8,
@@
and the result follows by Lemma B.12. ∎
Proposition B.14.
@@
$\begin{split}{{\left\|\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right\|}}_{F}^{2}\leqslant 2\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\Delta_{{\boldsymbol{\hat{A}}}}^{\top}\left(I-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}E\right)\,.\end{split}$ (25)
@@
By Lemma B.8,
$\begin{split}\mathcal{L}({\boldsymbol{A}})&={{\left\|\left(M_{{\boldsymbol{A}}}-I\right)L_{\star}E\right\|}}_{2}^{2}={{\left\|\left[\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}}\right)L_{\star}-I\right]E\right\|}}_{F}^{2}\\ &={{\left\|\left[\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}+\left(M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}-M_{{\boldsymbol{A}}^{\star}}\right)L_{\star}-I\right]E\right\|}}_{F}^{2}\\ &={{\left\|\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}E\right\|}}_{F}^{2}+{{\left\|\left[\left(M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}-M_{{\boldsymbol{A}}^{\star}}\right)L_{\star}-I\right]E\right\|}}_{F}^{2}\\ &\quad\quad+2\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)^{\top}\left[\left(M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}-M_{{\boldsymbol{A}}^{\star}}\right)L_{\star}-I\right]E\right)\\ &={{\left\|\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}E\right\|}}_{F}^{2}+{{\left\|\left(M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}-I\right)L_{\star}E\right\|}}_{F}^{2}+2\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)^{\top}\left(M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}-I\right)L_{\star}E\right)\,.\end{split}$
Then, for ${\boldsymbol{\hat{A}}}$ that satisfy $\mathcal{L}({\boldsymbol{\hat{A}}})\leqslant\mathcal{L}({\boldsymbol{A}}^{\star}_{p^{\prime}})$, we have
$\begin{split}\mathcal{L}({\boldsymbol{\hat{A}}})-\mathcal{L}({\boldsymbol{A}}^{\star}_{p^{\prime}})&={{\left\|\left(M_{{\boldsymbol{\hat{A}}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}E\right\|}}_{F}^{2}+2\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\left(M_{{\boldsymbol{\hat{A}}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)^{\top}\left(M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}-I\right)L_{\star}E\right)\leqslant 0\,,\end{split}$
which implies the desired result. ∎
Corollary B.15.
Observe that for $p^{\prime}=p$, Proposition B.14 reads
${{\left\|\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right\|}}_{2}^{2}\leqslant 2\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right)\,.$

### v2

@@ B.4 Covering numbers
Since for any matrix $M$, $\|M\|_{\rm op}\leqslant\|M\|_{F}$, it is sufficient to give a covering number for the Frobenius norm. Let $\mathcal{M}_{r}^{d\times d}(D)$ be the set of low-rank matrices of rank $r$ and Frobenius norm bounded by $D$:
$\mathcal{M}_{r}^{d\times d}(D)={{\left\{M\in\mathbb{R}^{d\times d}\mid\|M\|_{F}\leqslant D,\ {\rm rank}(M_{i})\leqslant r\right\}}}\,.$
By Lemma B.3, there exists an $\dfrac{\epsilon}{p^{\prime}}$-net $\mathcal{S}$ of $\mathcal{M}_{r}^{d\times d}(D)$ for Frobenius norm such that
$|\mathcal{S}|\leqslant\left(\dfrac{9Dp^{\prime}}{\epsilon}\right)^{(2d+1)r}\leqslant\exp\left(9dr\ln\dfrac{Dp^{\prime}}{\epsilon}\right)\,,$
as any $\dfrac{\epsilon}{Dp^{\prime}}$-net of $\mathcal{M}_{r}^{d\times d}(1)$ gives an $\dfrac{\epsilon}{p^{\prime}}$-net of $\mathcal{M}_{r}^{d\times d}(D)$.
Observe that the set $\mathcal{M}_{r,p^{\prime}}(D)$ is a subset of
$\begin{split}\mathcal{U}&=\left(\mathcal{M}_{r}^{d\times d}(D)\right)^{p^{\prime}}\times{{\left\{0_{d\times d}\right\}}}^{p-p^{\prime}}\,.\end{split}$
Then, $(\mathcal{S})^{p^{\prime}}$ gives an $\epsilon$-net of $\mathcal{U}$ and hence of $\mathcal{M}_{r,p^{\prime}}(D)$. ∎
B.5 Proof of Proposition 3.1
Proof.
Using Lemma C.4, we have ${{\left\|M_{{\boldsymbol{A}}^{\star}}\right\|}}_{\rm op}\leqslant\sqrt{p}{{\left\|{\boldsymbol{A}}^{\star}\right\|}}_{\rm op}$, directly leading to (ii).
@@
Appendix C Proof of Theorems 4.1, 4.2 and 4.3
@@
Proposition C.1.
With the definitions of Definition C.1, we have the following properties:
@@
$\begin{split}\mathcal{A}_{r,p^{\prime}}(D)&={{\left\{{\boldsymbol{A}}\in\mathbb{R}^{d\times pd}\ \Big{|}\ \|\Delta_{{\boldsymbol{A}}}\|_{\rm op}\leqslant D,\ {\rm rank}(A_{i})\leqslant r,\ A_{p^{\prime}+1}=\cdots=A_{p}=0\right\}}}\,,\\ \mathcal{S}_{r,p^{\prime}}(C,D)&={{\left\{{\boldsymbol{A}}\in\mathcal{A}(D)\ \big{|}\ \|{\boldsymbol{A}}-{\boldsymbol{A}}^{\star}_{p^{\prime}}\|_{F}^{2}\leqslant CD^{2}\eta^{2}\tau\dfrac{p^{\prime}dr}{N(T-p^{\prime})}\right\}}}\,,\end{split}$
@@
and $\tau$ is the following logarithmic term:
$\tau=\left(1+\ln\dfrac{p^{\prime 2}dNT^{2}}{T-p^{\prime}}\right)^{3}\left(1+\ln\kappa\right)\,.$
@@
$\mathcal{G}_{r,p^{\prime}}(C,D)={{\left\{{\boldsymbol{A}}\in\mathcal{A}_{r,p}(D)\ \Big{|}\ \dfrac{\|\Delta_{{\boldsymbol{A}}}\|_{F}^{2}}{\|\Delta_{{\boldsymbol{A}}}\|_{\rm op}^{2}}\leqslant C\eta^{2}\tau\frac{p^{\prime}dr}{N}\right\}}}\,.$
We set $\mathcal{A}_{r,p^{\prime}}=\mathcal{A}_{r,p^{\prime}}(\infty)$ and $\mathcal{G}(C)_{r,p^{\prime}}=\mathcal{G}(C,\infty)$. Lastly, we drop the subscript $r,p^{\prime}$ when it is clear from the context.
C.1 Main results
In this subsection, we state Theorem C.1 that generalizes the statements in Theorems 4.1, 4.2 and 4.3. We give a proof that reduces Theorem C.1 to a uniform concentration result in Theorem C.2. The proof of Theorem C.2 is deferred to Section C.5. Next, in Remark C.1, we discuss the factor $\eta$ that appears in our results and the conditions under which our misspecification bounds are tight. Lastly, Corollary C.1 removes the constraint on the diameter of the search set which allows us to treat the OLS estimator as a special case of the main theorem.
Theorem C.1.
Let Assumptions 3.1 and 3.2 hold. Furthermore, let Assumption 3.3 for $r<d$ and Assumption 3.5 for $p^{\prime}<p$ hold. Consider the following constrained empirical risk minimizer:
@@ C.1 Main results
Then, for any $0<\delta<e^{-1}$, there exists a constant $C(\delta)$ such that
@@
By Corollary C.3, $\mathcal{G}(C,D)\subset\mathcal{S}(C,D)$ and thus, for any random choice of ${\boldsymbol{A}}$,
($\begin{split}\mathbb{P}\left({{\left\{{\boldsymbol{A}}\in\mathcal{S}(C,D)\right\}}}\right)&\geqslant\mathbb{P}\left({{\left\{{\boldsymbol{A}}\in\mathcal{G}(C,D)\right\}}}\right)=1-\mathbb{P}\left({{\left\{{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)\right\}}}\right)\,.\\ \end{split}$
For the choice of ${\boldsymbol{\hat{A}}}$, $\mathbb{P}\left(\mathcal{E}_{{\boldsymbol{\hat{A}}}}\right)=1$ by Corollary C.4 and
($\mathbb{P}\left({{\left\{{\boldsymbol{\hat{A}}}\in\mathcal{S}(C,D)\right\}}}\right)\geqslant 1-\mathbb{P}\left({{\left\{{\boldsymbol{\hat{A}}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)\right\}}}\mid\mathcal{E}_{\boldsymbol{\hat{A}}}\right)\,.$
@@
($\begin{split}\mathbb{P}\left({{\left\{{\boldsymbol{\hat{A}}}\in\mathcal{S}(C,D)\right\}}}\right)&\geqslant 1-\mathbb{P}\left(\mathcal{E}_{\boldsymbol{\hat{A}}}\mid{{\left\{{\boldsymbol{\hat{A}}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)\right\}}}\right)\mathbb{P}\left({{\left\{{\boldsymbol{\hat{A}}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)\right\}}}\right)\\ &\geqslant 1-\mathbb{P}\left(\mathcal{E}_{\boldsymbol{\hat{A}}}\mid{{\left\{{\boldsymbol{\hat{A}}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)\right\}}}\right)\,.\end{split}$
Then, the proof is complete by applying Theorem C.2 to the right-hand side. ∎
Theorem C.2.
Let all the assumptions of Theorem C.1 hold. Then, for any $0<\delta<e^{-1}$, there exists a constant $C(\delta)$ such that
@@
Remark C.1.
For strictly stable systems with $\|M_{{\boldsymbol{A}}^{\star}}\|_{\rm op}<1$ and $\|M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\|_{\rm op}<1$, the factor $\eta$ is controlled by Corollary C.2. However, for marginally stable systems or explosive systems, there is no a prior good upper bound on $\eta$, implying that the misspecification results only apply to strictly stable systems without any further assumptions.
Corollary C.1.
Let all the assumptions of Theorem C.1 hold and suppose that $NT$ verifies the following condition:
$N(T-p^{\prime})\geqslant C(\delta)\eta^{2}\tau p^{\prime 2}dr\,,$ (20)
where $0<\delta<e^{-1}$ is a constant and $C(\delta)$ is given by Theorem C.1 Then, the OLS estimator
${\boldsymbol{\hat{A}}}_{\rm{OLS}}=\operatorname{argmin}_{{\boldsymbol{A}}\in\mathcal{A}(\infty)}\mathcal{L}({\boldsymbol{A}})\,,$
satisfies the same concentration result as in Theorem C.1:
$\mathbb{P}\left({\boldsymbol{\hat{A}}}_{\rm{OLS}}\in\mathcal{S}(C(\delta),D)\right)\geqslant 1-\delta\,.$
@@
$\mathcal{A}(D)^{\mathrm{o}}\subset\mathcal{S}(C(\delta),D)\,,$
i.e., the interior of $\mathcal{A}(D)$ contains $\mathcal{S}(C(\delta),D)$. We have the following relation:
$\mathcal{A}(\infty)={{\left\{{\boldsymbol{A}}^{\prime}=\alpha{\boldsymbol{A}}\mid{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{S}(C,D),\alpha\geqslant 1\in\mathbb{R}\right\}}}\,.$
Then, by Theorem C.2, we have
$\begin{split}&\mathbb{P}\left(\exists{\boldsymbol{A}}\in\mathcal{A}(\infty)\setminus\mathcal{S}(C(\delta),D):\|\Delta_{\boldsymbol{A}}L_{\star}E\|_{F}^{2}\leqslant 2\eta\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)\right)\\ &\quad\quad\quad=\mathbb{P}\left(\exists\alpha\geqslant 1\in\mathbb{R},{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{S}(C(\delta),D):\|\Delta_{\alpha{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}\leqslant 2\eta\operatorname{Tr}(E^{\top}\Delta_{\alpha{\boldsymbol{A}}}L_{\star}E)\right)\\ &\quad\quad\quad=\mathbb{P}\left(\exists{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{S}(C(\delta),D):\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}\leqslant 2\eta\operatorname{Tr}(E^{\top}\Delta_{{\boldsymbol{A}}}L_{\star}E)\right)\leqslant\delta\,,\end{split}$
@@
Thus, the result is complete by applying the same argument as in Theorem C.1 where $\mathcal{A}(D)$ is replaced by $\mathcal{A}(\infty)$.
We only need to provide a $D$ such that $\mathcal{A}(D)^{\mathrm{o}}\subset\mathcal{S}(C(\delta),D)$. For any ${\boldsymbol{A}}\in\mathcal{S}(C(\delta),D)$, we have
$\|\Delta_{{\boldsymbol{A}}}\|_{\rm op}^{2}\leqslant p^{\prime}\|{\boldsymbol{A}}\|_{\rm op}^{2}\leqslant p^{\prime}\|{\boldsymbol{A}}\|_{F}^{2}\leqslant C(\delta)D^{2}\eta^{2}\tau\dfrac{p^{\prime 2}dr}{N(T-p^{\prime})}\,,$
from Lemma C.4. Therefore, we need to find a $D$ such that
$D^{2}\geqslant C(\delta)D^{2}\eta^{2}\tau\dfrac{p^{\prime 2}dr}{N(T-p^{\prime})}\,,$
which equivalent to the condition in Equation 20. ∎
C.2 Technical lemmas
In this subsection, we present simple technical results on $L_{\star},M_{{\boldsymbol{A}}}$ and $\Delta_{{\boldsymbol{A}}}$ that are used in the proof of Theorem C.1.
Lemma C.1.
@@ C.2 Technical lemmas
By Weyl’s inequality for singular values on the identity $L_{\star}=M_{{\boldsymbol{A}}^{\star}}L_{\star}+I$ from Lemma C.1,
@@
This implies the desired inequalities for $\|L_{\star}\|_{\rm op}$. For the lower bound on minimal singular value, use Lemma C.1,
@@
Applying Lemma C.2,
$\eta\leqslant\left\|M_{{\boldsymbol{A}}^{\star}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right\|_{\rm op}\|L_{\star}\|_{\rm op}\leqslant\frac{\left\|M_{{\boldsymbol{A}}^{\star}}\right\|_{\rm op}+\left\|M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right\|_{\rm op}}{1-\|M_{{\boldsymbol{A}}^{\star}}\|_{\rm op}}\leqslant\dfrac{2}{1-\|M_{{\boldsymbol{A}}^{\star}}\|_{\rm op}}\,.$
Lemma C.3.
@@
By Weyl’s inequality for singular values on the identity $L_{\star}=I+M_{{\boldsymbol{A}}^{\star}}+\cdot+M_{{\boldsymbol{A}}^{\star}}^{T-1}$ from Lemma C.1,
@@
For the lower bound on minimal singular value, use Lemma C.1,
@@
and the result follows by Lemma C.4. ∎
Proposition C.2.
@@
${{\left\|\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right\|}}_{F}^{2}\leqslant 2\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\Delta_{{\boldsymbol{\hat{A}}}}^{\top}\left(I-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}E\right)\,.$
@@
By Lemma C.1,
$\begin{split}NT\mathcal{L}({\boldsymbol{A}})&={{\left\|\left(M_{{\boldsymbol{A}}}-I\right)L_{\star}E\right\|}}_{2}^{2}={{\left\|\left[\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}}\right)L_{\star}-I\right]E\right\|}}_{F}^{2}\\ &={{\left\|\left[\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}+\left(M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}-M_{{\boldsymbol{A}}^{\star}}\right)L_{\star}-I\right]E\right\|}}_{F}^{2}\\ &={{\left\|\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}E\right\|}}_{F}^{2}+{{\left\|\left[\left(M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}-M_{{\boldsymbol{A}}^{\star}}\right)L_{\star}-I\right]E\right\|}}_{F}^{2}\\ &\quad\quad+2\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)^{\top}\left[\left(M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}-M_{{\boldsymbol{A}}^{\star}}\right)L_{\star}-I\right]E\right)\\ &={{\left\|\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}E\right\|}}_{F}^{2}+{{\left\|\left(M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}-I\right)L_{\star}E\right\|}}_{F}^{2}\\ &\quad\quad+2\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)^{\top}\left(M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}-I\right)L_{\star}E\right)\,.\end{split}$
Then, we have
$\begin{split}NT\left(\mathcal{L}({\boldsymbol{A}})-\mathcal{L}({\boldsymbol{A}}^{\star}_{p^{\prime}})\right)&={{\left\|\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}E\right\|}}_{F}^{2}\\ &\quad\quad+2\operatorname{Tr}\left(E^{\top}L_{\star}^{\top}\left(M_{{\boldsymbol{A}}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)^{\top}\left(M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}-I\right)L_{\star}E\right)\,,\end{split}$ (22)
which implies the desired result for any ${\boldsymbol{A}}$ that satisfies $\mathcal{L}({\boldsymbol{A}})\leqslant\mathcal{L}({\boldsymbol{A}}^{\star}_{p^{\prime}})$. ∎
Corollary C.4.
Observe that for $p^{\prime}=p$, Proposition C.2 reads
${{\left\|\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right\|}}_{F}^{2}\leqslant 2\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right)\,.$